OpenStudy (anonymous):

For the function 32/(x+3), find (g^-1∘g)(5).

OpenStudy (anonymous):

First fidn g^-1. To do that, write y= the function, switch y and x and solve for y. y=32/(x+3) x=32/(y+3) When you solve for y that function will be your g^-1. Then it says compute g(5), meaning plug in 5 into your original equation. Then Compute g^-1(g(5)), meaning plug in the value you got from g(5) into your inverse function.